OpenStudy (anonymous):

the sum of two numbers is 16 and the sum of their squares is 146. find the numbers

OpenStudy (cwrw238):

let the 2 numbers be x and y x + y = 16 x^2 + y^2 = 146 from first equation y = 16 - x substitiue for y in second equation; x^2 + (16 -x )^2 = 146 solve this for x

OpenStudy (cwrw238):

can you solve that equation mike12?

OpenStudy (anonymous):

no still working on it...dont really understand it